### v1

# New Coevolution Dynamic as an Optimization Strategy in Group Problem Solving

Abstract
In society, it is common to face problems that require collaboration with other people, from everyday challenges to complex tasks, such as group projects at work. In this context, the search for more effective problem-solving strategies becomes a topic of great interest. This paper presents a new dynamic that integrates coevolution into the NK model together with an Erdös-Rényi random network, allowing more than one neighbor update. We explore how this coevolution can be employed as an optimization strategy for solving group problems. In recent coevolution models, only one agent is removed and another is added to the neighborhood of influence. Here, we allow $L$ agents to be changed, which had a large impact on the system performance. In the analysis of the results, we consider rewiring as a way for the target agent to obtain information from individuals or groups that were not part of its neighborhood. Our simulations demonstrate that coevolution can produces gain on the computational cost.
Keywords Coevolution, Group Problem Solving, NK model, Random Network
1 Introduction
Social interactions are inherently complex such that in recent years numerous models of social behavior using Network Science have emerged to explore this complexity [1]. Each new model provides a fresh perspective on human interaction [2]. Presently, a prominent aspect under exploration within complex network models is coevolution. In many complex systems the dynamic processes that occur in a network are interconnected with changes in the network topology [3, 4]. The rationale behind this focus is that co-evolutionary, or adaptive, network models offer a more accurate portrayal of real world systems compared to static or evolving networks. Empirical networks display both network dynamics (the evolution of network topology) and node state dynamics. Therefore, coevolution can be conceptualized as two simultaneous processes – link rewiring and state dynamics – mutually influencing each other [5, 6, 7]. A network is composed of a set of nodes (where each node represents an element of the system) connected by links (where each link represents the interaction between the nodes). Despite being simple, this structure is capable of encoding information present in a wide variety of natural, artificial and social interaction systems [8, 9].
When considering cooperative processes, it is acknowledged that analyzing a group performance in problem-solving can be intricate [10]. Initially, one might assume that a larger number of individuals working towards the same objective would expedite reaching a solution. However, interpersonal interactions (such as everyone engaging in discussion) can lead to suboptimal outcomes, deviating from the desired solution. Different types of networks are found in the literature (square network, RGG, Erdös and Rényi, among others) and the choice of each one depends on the type of problem one wants to address. This is because the network determines the model communication pattern: who interacts with whom. The model determines what happens when someone interacts with someone.
Given the way our society is structured, human cooperation is crucial for social and technological development. We can understand this cooperation as interactions between peers and coordinated behavior in larger groups [11]. From a social perspective, each node in the network can represent an individual or a group of individuals, and each link represents the connection with people or groups that have a connection with the target agent. In this study, we consider rewiring as a way for the target agent to obtain information from individuals or groups that were not previously part of its neighborhood.
In this work we present a new dynamics with coevolution considering an Erdös-Rényi random network and we use the NK model [12]. Usually in coevolution models only one agent is removed and another one is added to the influence neighborhood. Here we allow $L$ agents to be altered, which had great impact on the performence of the system. This work is organized as follows: In Section II we present a brief discussion about the NK model and the construction of reliefs. In Section III we briefly discuss the Erdös-Rényi random network and present the proposed new co-evolution dynamics. Section IV is dedicated to presenting the results obtained. Section V presents our final considerations.
2 Model
2.1 NK Fitness Landscapes
[…]
For $K=0$ the function $f(\phi)$ has only one maximum $f(\phi_{M})=f_{M}$, the global one. Therefore the difference $f(\phi_{M})-f(\phi)$ is proportional to the Hamming distance $d_{H}(\phi_{M},\phi)$ between the vectors. The Hamming distance gives the number of different bits between the two vectors. For example $d_{H}(\phi_{1},\phi_{2})=1$ for $\phi_{1}=(010101)$ and $\phi_{2}=(000101)$. With $K>0$ the problem becomes more complicated because the function $f$ has additional local maxima beside the global one. Thus the difference $f(\phi_{m})-f(\phi)$ is no longer proportional to $d_{H}(\phi_{m},\phi)$. The average number of local maxima increases with $K$ until the limiting case of $K=N-1$, in which the landscape becomes completely random, known as the Random Energy model [12].
The dynamics of the temporal evolution begins at $t=0$ with the selection of an agent at random for a string update. We shall call this selected agent as the target one. This update can happen by two different ways: copy or mutation. The mutation happens when the target agent has no neighbors or when it is the model of its influence neighborhood (which is the set of neighbors agents of the target). Been the model means that the target agent has the highest fitness of his neighborhood. On the other hand, the copy will happen when the target has neighbors and it is not the model. Both processes change the string of the target agent. The mutation is achieved by a flip of a random bit of the string. The copy process involves more steps: the different bits bewteen the target and the model strings are obtained. Then, one of them is randomly selected and flipped on the target string. As a result, the target string gets more similar to the model string.
The efficiency of coevolution is measured by the total number of iterations $t_{g}$ required by the first agent who finds the global maximum. A smaller $t_{g}$ means more efficient search: the goal was achieved in less time. One iteration refers to $N$ updates of $N$ randomly selected agents. The normalized computational cost is then defined as $C\equiv Mt_{g}/2^{N}$, so it can be compared between systems of different size $M$.
The independent search consider $M$ agents and the updates are only mutation. This case is used as an reference and has an analytic result [14]:
$\left\langle C\right\rangle=\frac{M}{2^{N}\left[1-\left(\lambda_{N}\right)^{M}\right]},$ (1)
where $\lambda_{N}$ depends on the value of $N$. From the definition of $C$ and Eq. 1, the expected number of trials for a group of $M$ independente agents to find the global maximum is:
$\left\langle t_{g}\right\rangle=\frac{1}{\left[1-\left(\lambda_{N}\right)^{M}\right]}.$ (2)
2.2 Random Network
A network or a graph comprises a collection of vertices interconnected by a set of edges [15]. Each edge, or connection, links a pair of vertices with the number of vertices defining the network size and the set of edges determining its connectivity. Vertices that share an edge are termed neighbors. Throughout this work the terms vertices, agents, and individuals have the same meaning.
The degree $k_{i}$ of a vertex refers to the number of connections it establishes with other vertices while the network connectivity is characterized by the distribution $P(k)$. Various types of networks are used for solutions of physical and biological problems. In this work we will use the Erdös-Rényi (ER) network [16, 17, 18, 19] which has the following definition: consider a set of $M$ vertices and a parameter $\beta$ such that $1.00.0<\beta<$ 1.0, the probability of any two vertices being connected is $\beta$. The average degree $\mu$ of the network is defined as $\mu=(1/M)\sum_{i}^{M}k_{i}$ and in the case of ER network it is $\mu=\beta(N-1)$ [20]. So we use $\mu$ as control parameter of the network instead of $\beta$.
2.3 Coevolution
The general idea of coevolution is to apply a rewiring before each string update: removing a random neighbor and adding a new one to the target agent [5]. However, we can let the number of destroyed and created connections be greater than one. Therefore, before each string update we remove $L<k_{i}$ random neighbors and add $L$ new random neighbors to the target agent. The parameter $L$ is an input to our model, it is the control parameter of the coevolution process.
The rewiring itself is the temporal evolution of the network. To be considered as coevolution, this network evolution must be coupled to the evolution of the agent string. The constant rewiring (always happening) is not coevolution, although is a network evolution. The no coevolution case is $L=0$. We have tested several different mechanisms to implement this coupling and found one which gave interesting results. The coevolution happens with a probability given by $r^{q}$ where:
$r=1-\frac{\phi_{t}-\phi_{s}}{\phi_{m}-\phi_{s}},$
where $\phi_{t}$ is the fitness of the target agent, $\phi_{s}$ is the smallest fitness and $\phi_{m}$ is the maximum one (fitness of the model agent) on the influence neighborhood. With this definition, if the target is the model agent we have $r=0.0$. If it has the lowest fitness $\phi_{s}$ we have $r=1.0$. Otherwise we have $1.00.0<r<$ 1.0. The coevolution ($q>0$) is a selective rewiring, in contrast with the constant rewiring case when $q=0$. When the coevolution happens, $L$ neighbors of the target agent are removed and new $L$ random ones are created. The parameter $q$ is the non-linear one: $q=1$ is the linear case, $q=0$ means constant rewiring (always happening) and otherwise is the non-linear case. This parameter $q$ is added to the rewiring probability definition to allow studies of possible non-linear effects, as observed by for example by Raducha et al [5].
3 Results
A schematic illustration for this new dynamic is presented in Figure 1. The steps of one iteration are:
  • At time $t=0$ we initialize the network and randomly select a node to serve as the target agent.
  • The rewiring happens with probability of $r^{q}$.
  • The string update can be either a mutation or a copy. Mutation arises when the target agent is the model one. Otherwise copying takes place with probability $p$ and mutation with probability $1-p$.
$N$ iteractions are one Monte Carlo Step (MCS). The simulation concludes when the target agent achieves the highest global fitness after $t_{g}$ MCS, which is one time evolution. The computational cost $\left\langle C\right\rangle$ is averaged over $S$ independent time evolutions, which one with a random sample of the network.
Figure 1: Schematic illustration of the coevolution dynamics used. The initial network is the left bottom one and the final network is the center top one. The colors represent different components. In the initial network the target node is enclosed within a black dashed rectangle. $r_{1}$ and $r_{2}$ are random numbers from a uniform distribution between 0 and 1. The parameter $q$ is the non-linear one: $q=1$ is the linear case, $q=0$ means constant rewiring (always happening) and otherwise is the non-linear case.
In this work we explore two values for the $K$ parameter: $K=0$ is the easy case and $K=4$ is the complicated one. We study the linear case $q=1$ except when it is indicated otherwise. We also fixed $N=12$ and the copy probability at $p=0.5$. At this value of $N$ we have $\lambda_{12}\approx 0.99978$ and $t_{1}=\left\langle t_{g}\right\rangle\approx 4545$ for a system with only one agent (from Eq. 2). Since $(\lambda_{12})^{M}\approx e^{-M(1-\lambda_{12})}$, we can write the approximated results [12]:
$\left\langle C\right\rangle\approx\left\{\begin{array}[]{rl}t_{1}/2^{N},&\text{ for }M<<t_{1}\\ M/2^{N},&\text{ for }M>>t_{1}\end{array}\right.$
At $N=12$, the limit value for low $M$ is $\left\langle C\right\rangle=t_{1}/2^{12}\approx 1.110$. We compare our results with 3 others data sets:
  • Independent Search (IS): when each agent independently searches for the global maximum. In this case there is no copy of the model strings, only mutation. It can be achieve by setting $p=0$ or using a network with $k_{i}=0$ for all agents $i$. This result will be displayed as a green solid line on the results.
  • Full Connected Network (FCN): when all agents are connected to all other agents: $k_{i}=M-1$ for all agents. This result will be displayed as a black dashed line on the results.
  • Reference Case (RC): dynamics with no coevolution, equivalent to $L=0$.
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [21], Matplotlib [22], NetworkX [23] and Pandas [24]) and Gnuplot. All codes and data are available upon request.
3.1 Reference case
Firstly we present the results with no coevolution: the Reference Case (RC). At $K=0$ (Figure 2(a)) the best result (lower $\left\langle C\right\rangle$) is given by the full connected network (FCN) and the worst by the independent search (IS). The increasing average degree $\mu$ interpolates from IS to the FCN case. As there is only one maximum (the global one) it becomes easier to increase the fitness with more neighbors to copy. However, in terms of system size $M$ there is one optimal value roughly between 15 and 20 which returns $\left\langle C\right\rangle\sim 0.007$. So, if there is too few neighbors (smaller $M$) or too much ones (larger $M$) the agents spend more time looking for the global maximum. At large $M\gtrsim 10^{4}$ the system behaviors as the limit case $M/2^{N}$.
Figure 2: Computational cost $\left\langle C\right\rangle$ as function of $M$ for different average degrees $\mu$ of the ERN with no coevolution. Parameter: $q=1$ and $S=5\times 10^{4}$ samples. The green line is the IS (Eq. 1), black dashed line is the FCN result and the pink line is the approximation for large $M$. Purple, blue, green, orange and red circles correspond to an average degree of $\mu=1$, 5, 10, 20 and 100, respectively. (a) $K=0$. (b) $K=4$.
When $K=4$ (Figure 2(b)) the scenario is drastically different. As before there is an optimal value of $M$ roughly between 10 and 15. However there is a significant increase on the cost $\left\langle C\right\rangle$ at $M\gtrsim 20$ when $\mu$ is increased towards the FCN. As there are local maxima in this landscape, an agent can be trapped in one of these making it longer to find the global maxima. The results on these two scenario resemble the similar case on the random geometric graph [25].
3.2 Dynamics with coevolution
In order to better enhance the effect of coevolution on the cost, we restricted the values of the average degree for the analysis: we have used $\mu=1.0$ and $\mu=10.0$ for the cases $K=0$ and $K=4$, respectively. We begin our study with the linear case: $q=1$. Considering the easy landscape with $K=0$ (Fig. 3(a)), the rewiring brings benefit to the cost except when $L=1$. As there is only one maximum, any other agent with a higher fitness is worth copying. And this process is facilitated by the selective rewiring of the coevolution. Interestingly, the case $L=1$ increases the cost. The difficult case $K=4$ (Fig. 3(b)) has some similarities: the cost increases at $L=1$ and then decreases up to $L=5$. At $L=6$ there is a significant increase only at $M\sim 60$.
Figure 3: Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of $L$. $L=0$ means no coevolution. Parameters: $q=1$, $S=10^{5}$ samples and $\mu=1$. The green line is the IS (Eq. 1) and the black dashed line is the FCN result. (a) $K=0$. (b) $K=4$.
At higher average degree $\mu$ the results are drastically different. The ease case $K=0$ (Fig. 4(a)) shows influence of $L$ only on small network sizes around $M\lesssim 15$. At $K=4$ the results become more interesting. At small networks coevolution continues to present a higher computational cost. However, at medium-sized networks ($20\leq M\leq 120$), we see a significant improvement in computational cost. $\left\langle C\right\rangle$ presents a maximum at $M=67$ for $L=0$ (Fig. 4(b)) and a minimum at $L=3$.
Figure 4: Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of the number of modifications $L$. Parameters: and $S=10^{5}$ and $\mu=10$. The green line is the IS (Eq. 1) and the black dashed line is the FCN result. (a) $K=0$. (b) $K=4$.
The $\left\langle C\right\rangle$ minimum created by $L$ has a larger sensitivity with $M$ than with $\mu$. As can be seen on Fig. 5(a), the cost has a pronounced minimum between $L=2$ and 3 considering different values of $M$. However when $M$ if fixed and $\mu$ is varied the minimum is not so evident (Fig. 5(b)), and even not present at some values.
Figure 5: Results of $\left\langle C\right\rangle$ vs. $L$. Parameters: $K=4$, $q=0$ and $S=30000$. (a) Different values of $M$ and $\mu=10$. (b) Different values of $\mu$ and $M=67$.
To better visualize the behavior of the $\left\langle C\right\rangle$ minimum at $M=67$, Fig. 6(a) displays a color map on the space $L,\mu$. However, a more interesting measure is the difference between the $L=3$ and $L=0$ values, defined as:
$\Delta(L,\mu)=\left\langle C\right\rangle(L,\mu)-\left\langle C\right\rangle(0,\mu).$ (3)
This quantity gives exactly the improvement (given by $\Delta<0$) on the cost $\left\langle C\right\rangle$ created by the coevolution, as can be seen on Fig. 6(b). The average degree as a great impact on this gain: at $L=6$ it goes from destroying at $\mu=6$ to improving the performance at $\mu=12$. This result shows the tunable coevolution described here is a real tool to increase perfomance on the NK model.
Figure 6: Results of $\left\langle C\right\rangle$ in the $(L,\mu)$ space. Parameters: $K=4$, $q=1$, $M=67$ and $S=10^{5}$. (a) $\log\left\langle C\right\rangle$. (b) $\Delta$ defined on Eq. 3. Positive values are indicated in red and negative values in blue.
3.3 Other dynamics
To verify the impact of the coevolution dynamics, we compare it with the constant rewiring one: $q=0$. Considering the curve $L=1$ from Fig. 4(b), the constant rewiring case gives a lower cost, as can be seen on Fig. 7(a). Focusing on the peak at $M\sim 67$, the results shows a slightly better performance (lower cost) with $L=1$ for the constant rewiring case, while negligible differences when $L=3$. Figures 7(b) and 7(c) show these same results for different average degrees $\mu$. One can see that the improvement from coevolution (or the constant rewiring) case is not present.
Figure 7: Computational cost ($\left\langle C\right\rangle$) as a function of network size $M$ for different values of the average degree. The green line is the IS (Eq. 1), the black dashed line is the FCN result and RC is the reference case (equivalent to $L=0$). In (a) $\mu=10$, (b) $\mu=5$ and (c) $\mu=15$.
Lastly we studied the impact of non-linearity on the computational cost, which happens when $q$ is different from 0 or 1. In this work non-linearity yield no significant outcome, as can be shown on Fig. 8. The case $q=0.5$ only showed slightly better performance than the linear case $q=1$ around $M=67$. Other than this region there were no significant differences.
Figure 8: Computational cost $\left\langle C\right\rangle$ vs. $M$ for different values of the non-linearity parameter $q$. Parameters: $\mu=10$, $S=3000$ and $L=3$. The violet circles correspond to $q=0.5$, the orange circles to $q=1.5$, and the red circles to $q=2$. The reference case RC ($L=0$ and $q=1$) is depicted in gray.
4 Conclusions
In this work, we investigate the impact of coevolution on the NK model within the context of an Erdös-Rényi network. Given that the NK model operates as a fitness model, our goal was to evaluate whether a tunable coevolution yields any gain in computational cost $C$. To this end, we introduce a new dynamic that incorporates the effect of rewiring by changing the neighborhood of the target agent. Notably, this study represents the first exploration of coevolution within the NK model using an ER network, along with the introduction of a new parameter linked to rewiring: the tunable parameter $L$. This parameter allows for rewiring interactions not limited to individual pairs of agents. Our simulations showed that the proposed dynamics can return a reduction in cost on the system. We also investigated the scenario with constant rewiring and have observed that the results are similar to the coevolution case. More studies can be needed to map the regions with improved perfomance for larger $L$. It can be also interesting to study the impact of this tunable coevolution on the phase diagram of with absorbing phase models.
Acknowledgements
This work received financial support from the brazillian agencies CNPq (project 405508/2021-2) and FAPEG (project no. 401425/2023-1). This research was also supported by LaMCAD/UFG.
Author contributions
Francis F. Franco: Conceptualization, Methodology, Software, Formal analysis, Investigation, Writing - Original Draft, Writing - Review e Editing, Paulo F. Gomes: Conceptualization, Methodology, Formal analysis, Investigation, Writing - Original Draft, Writing - Review, Editing, Supervision, Project administration. Authors declare that there are no competing interests.
Data Availability Statement
The model was implemented using Fortran and the analysis and graphics were developed using Python (packages: Numpy [21], Matplotlib [22], NetworkX [23] and Pandas [24]) and Gnuplot. All codes and data are available upon request.